OpenStudy (anonymous):

Which polynomial is a perfect square trinomial? 9x^2 − 12x + 4 36b^2 − 24b + 16 16x^2 − 24x − 9 4a^2 − 10a − 25

OpenStudy (anonymous):

Perfect square has the form \[(a \pm b)^2=a^2 \pm 2ab +b^2\] So you can see from the first equation that a² = 9x² This mean a = 3x You can also see that 4 = b² So b = 2 Now see if -12x = -2ab -2ab = 2*3x*3 = -12x This mean the first equation IS a perfect square. 9x² − 12x + 4 = (3x - 2)² So you have your answer already. But you could try this out for the other equations, too, for the sake of practice. :)

OpenStudy (anonymous):

Alright & thanks.. :) i thought it was the first one :)